OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

A binomial radical expression is of the form a + √b, where b is a positive rational but not a perfect square and a is rational. Its conjugate is a ‒ √b. It is of interest since their product is not a radical. To wit: (a + √b)(a ‒ √b) = a² ‒ a√b + a√b + (√b)² = a² + 0 + b² = a² + b² The middle terms of this FOIL application are opposites of each other and add to 0.

OpenStudy (anonymous):

thanks for your help on this one !

OpenStudy (anonymous):

do you think your able to help me on the problem i just showed you?

OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0.